Last year Marla purchased 100 shares of stocks for $8 per share. She paid a flat $75 to purchase the shares. Since making her purchase, she has received $200 in dividends. Marla is concerned that the stock price will fall below its current FMV of $7. Calculate her holding-period return if she sells today and pays a $75 commission.

Swarn bought 200 shares of a stock for $36 per share. He paid $245 in trading commissions. He received dividends in the amounts of $98, $156, and $300 over the last three years, respectively. Assuming that Swarn is in the 15% marginal tax bracket for capital gains and dividends, what is his holding period return if he sells all of the shares at $40 each with a $245 trading commission? What is Swarn's tax-adjusted holding period rate of return?
Answer #1

For Marla,

Number of shares purchased is 100 @ $8 per share

Initial Position = 100 * 8 = $800

Dividends Received = $200

Brokerage Paid = 75 + 75 = $150

Current Value of Stock = $7

Current Position for 100 shares @ $7 per share

Current Position = $700

Holding Period Return = Profit/Initial Investment

Holding Period Return = ((700 - 800) - 150 + 200)/800

Holding Period Return = -6.25%

For Swarn,

Initial Position is for 200 shares @ $36 per share

Initial Position = $7200

Brokerage Paid = 245 + 245 = $490

Dividends Received = 98 + 156 + 300 = $554

Current Position = 200 * 40 = $8000

Capital Gain = $800

Tax on Dividends and Capital Gain = 0.15(800 + 554)

Tax = $203.1

Holding Period Return = (800 + 554 -203.1 - 490)/7200

Holding Period Return = 9.18%